In the short-run production period, given the total cost of production as:                       

TC = wL + rK, where K is fixed

Prove that w/AP and SMC=w/MP 


1
Expert's answer
2022-01-06T09:08:50-0500

Solution:


The relation between average product and average cost:

Let y be the output produced by z1 units of input 1:

y = TP(z1).

Then we have

"AVC(y)=\frac{VC(y)}{y}=\frac{w_1z_1}{TP(z_1)}" ,

or

"AVC(y)=\frac{w_1}{\frac{TP(z_1)}{z_1}}" .

But

"\frac{TP(z_1)}{z_1}=AP(z_1)" ,

so that

"AVC(y)=\frac{w_1}{AP(z_1)}".


The relation between marginal product and marginal cost:


Take some amount z1 of input 1 and increase it a bit. The resulting increase in cost is

c = w1z1.

The change in the output is

y = MP(z1)z1.

Now, short-run marginal cost is the rate of change of variable cost: "\frac{c}{y}". So substituting we have


"SMC(y)=\frac{w_1}{MP(z_1)}" .